OpenStudy (anonymous):

How far from the base of a building must a 10ft ladder be placed so that it reaches 8ft up the wall?

OpenStudy (anonymous):

think of a 3 - 4 - 5 right triangle, then multiply everything by 2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

It would have to be 6 feet away from the base

OpenStudy (anonymous):

u could use the pythagorean theorem a2+b2=c2 (the 2s are squared)

OpenStudy (anonymous):

\[a^2+b^2=c^2\]